# Gravitational wave probes of particle dark matter: a review

### v3

@@ V.4.4 Ionization
Another transition arises when the binary system separation becomes comparable to the size of the cloud: an “ionization” of the boson cloud occurs as the cloud becomes unbounded from its host BH Takahashi et al. (2022); Baumann et al. (2022a, b). The energy from the binary sources this effect, and is significantly larger than that emitted via GWs; thus, the inspiral of the binary becomes driven by the boson cloud, instead of simply perturbed by it. Ionization will also tend to circularize the binary if it formed through dynamical capture, while leaving any orbital inclinations unaffected Tomaselli et al. (2023). It has been recently found, however, that as the resonances in the boson cloud evolve during the orbit, the boson cloud is unperturbed and ordinarily visible in LISA if the cloud and binary are counter rotating; otherwise, the cloud is destroyed because of the resonances, but a distinct mark is left on the binary. That is, the binary is forced to co-rotate with the boson cloud, and its eccentricity is driven to a particular value, which allows the possibilty of doing statistics with detected binary black hole systems in LISA to probe the boson cloud hypothesis Tomaselli et al. (2024).
@@ VI GW constraints on boson clouds
The extensive theoretical background on boson clouds around rotating BHs, coupled with the potential for them to be DM, have motivated GW probes of boson cloud/BH systems. Different GW interferometers probe vastly different mass regimes of ultralight dark matter boson clouds. While ground-based detectors are sensitive to annihilation signals from bosons of masses between $\sim 10^{-14}-10^{-12}$ eV Arvanitaki et al. (2015b), space-based detectors can probe a few orders of magnitude lower than that, $\sim 10^{-20}-10^{-16}$ eV Baumann et al. (2019a), and HFGW detectors could see systems with masses around $\sim 10^{-9}$ eV Aggarwal et al. (2021). At the moment, only ground-based GW interferometers exist, and so we focus on methods and search results from the most recent LIGO, Virgo and KAGRA observing runs.
VI.1 Constraints from the spins of detected binary black holes
We note that constraints on boson clouds can be inferred from the spin measurements of binary black hole systems Arvanitaki et al. (2015b, 2017); Brito et al. (2017); Baryakhtar et al. (2017); Cardoso et al. (2018); Ghosh and Sachdeva (2021). Recently, a Bayesian method has been developed by combining the information from each system Ng et al. (2021a), assuming a distribution for the initial BH spins and requiring that the BHs have had enough time to undergo superradiance before merging. Such a study has been applied to the second GW transient catalog (GWTC-2) Abbott et al. (2021b), and has resulted in essentially a joint posterior on the scalar boson mass and the spin distribution of BHs at their formation Ng et al. (2021b). The latter is required because BHs born with high spins that have been spun down by superradiance is partially degenerate with BHs simply born with low spins in a universe without superradiance. The study presented in Ng et al. (2021b) tentatively rules out scalar bosons in the mass range $[1.3,2.7]\times 10^{-13}$ eV.
@@ VI.2 Search methods
While the aforementioned constraints from measured BH spins can disfavor particular masses or even confirm the presence of an ultralight boson, they rely on assumptions about the spin distribution at formation, on whether the boson clouds have time to form (which could be impacted by eccentricity at formation Peters (1964); Wen and Gair (2005)), and are a function of the (strongest) BH mergers that we see with GWs. To that end, different methods to actually search for long-lived GWs from isolated BHs (or those present in a known x-ray binary system) have been developed and used in real searches in LIGO, Virgo and KAGRA data. Detecting GW emission from an isolated BH would provide smoking-gun evidence of the existence of an ultralight bosonic field around BHs.
VI.2.1 Excess power
One method to search for boson clouds around rotating BHs lightly models the boson cloud signal as monochromatic within the entire observation time that could come from anywhere in the sky, hence inducing a Doppler modulation. This method creates time-frequency maps and corrects each of them for different sky positions, leaving behind a monochromatic signal D’Antonio et al. (2018). It employs different coherence times $T_{\text{FFT}}$ for the analysis in order to be robust against possible theoretical uncertainties in the GW signal model. Shorter $T_{\text{FFT}}$, while less sensitive than longer ones, provide larger frequency bins that allow the signal frequency to wander more, and thus are more sensitive to deviations from the monochromatic signal model.
In LABEL:fig:tfmaps, we show the core stage of this method, which involves creating the time-frequency peakmap from raw time-domain strain data, correcting for the Doppler modulation from a particular sky location, and then integrating over time to collect all peaks at a given frequency to highlight the presence of a signal. Note that while we show the color to indicate the ratio of signal to noise power, the method does not actually use it: it only requires that a given point in the time-frequency plane is above a given threshold. When moving from LABEL:fig:afterdopp to LABEL:fig:pmproj, all time-frequency points above this threshold are given a weight of “1”, and the ones are summed over time in order to create this histogram.
@@ VI.2.2 Viterbi
Another method that is more robust against stochastic frequency variations relies on Hidden Markov Models (HMMs) to find arbitrary tracks in the time-frequency plane. The main idea is to model a time-frequency GW signal probabilistically as a Markov chain of transitions between “hidden” (unobservable) frequency states and use a detection statistic to relate “observed” frequency states with the hidden ones Quinn and Hannan (2001). In the framework for this method, the probability to jump to a new state at a given time only depends upon the previous hidden state. In total, there would be $N_{Q}^{N_{T}+1}$ possible paths through the hidden states, where $N_{Q}$ is the total number of hidden states and $N_{T}$ is the total number of times at which we have observed frequency states.The Viterbi algorithm provides a recursive, computationally efficient way of maximizing the probability that the hidden set of states is responsible for the observed sequence of states.
In LABEL:fig:sample-path, taken from Isi et al. (2019), we show the optimal path found by the Viterbi method for injected boson cloud signals in white noise that has an associated random walk of the frequency, which is meant to simulate unknown theoretical features of the signal. We can see that in both cases, with less and more variability of the signal, the Viterbi algorithm can find the appropriate track of the signal.
@@ VI.2.3 Cross-correlation and other methods
We have already discussed in Section III.1.1 cross correlating detector data to search for ultralight dark matter that could interact directly with GW interferometers; however, it is important to note that such a method could also be used to search for boson clouds around rotating BHs, both deterministically and for a stochastic GW background composed of the superposition of cosmological or astrophysical GW signals. Furthermore, virtually any method in the CW community could be tuned to look for such boson cloud systems because such methods already analyze almost monochromatic signals focusing on spin-downs but could also extend their parameter spaces to include spin-ups.
VI.3 Recent search results for scalar boson clouds
We can choose different BHs to point at depending on whether we are looking for scalar or vector clouds, or annihilation or transition signals. Current GW interferometer and algorithm sensitivities indicate that annihilation signals from scalar boson clouds could be observed for galactic and relatively young black holes Abbott et al. (2022d). Or, particular, known black hole systems, such as Cyg-X1, could be targeted for such clouds as well Sun et al. (2020). Finally, the stochastic GW background, composed of the superposition of signals from all of the annihilating boson cloud systems in the universe, can be probed as well through cross-correlation analyses. We go through each of these scenarios in the following subsections.
@@ VI.3.1 All-sky search constraints
The excess power method described in Section VI.2.1 has been used to analyze LIGO O3 data to search for the presence of annihilating scalar boson clouds around isolated BHs within our galaxy Abbott et al. (2022d). This search was very computationally expensive, since each sky position had to be analyzed individually. A separate search towards the galactic center was also performed, targeting both spinning down neutron stars and spinning up boson cloud systems Abbott et al. (2022e) No evidence for scalar boson cloud systems were found in either case, and thus upper limits were placed on the presence of boson clouds with different ages, spins, and distances away from earth in two different ways Palomba et al. (2019); Dergachev and Papa (2019); Abbott et al. (2022d):
@@
The first way is agnostic towards whether such BHs exist in the galaxy, while the second one contains some (uncertain) astrophysics regarding the properties of astrophysical BHs, but allows for some physical intuition into the reach of such searches.
@@
Figure 18: Taken from Abbott et al. (2022d). Using upper limits from the O3 LIGO search for GWs from annihilating boson cloud systems, exclusion region of particular boson mass /BH mass pairs are derived assuming different ages and distances (left: 1kpc; right: 15 kpc) for BHs with initial spins of $\chi_{i}=0.9$. This is the most model-agnostic way to probe boson clouds around rotating BHs using upper limits obtained from an all-sky search for CWs from isolated BHs.
Constraints from the second way are shown in Fig. 19, in which a Kroupa BH mass distribution $f(m)\propto M_{\rm BH}^{-2.3}$ Elbert et al. (2018) between $[5,~{}50]M_{\odot}$ and $[5,~{}100]M_{\odot}$ is assumed, as well as a uniform initial BH spin distribution between [0.2,0.9]. For fixed BH age, the maximum distance reach as a function of boson mass indicates that at least 5% of GWs from BH/boson cloud systems would have been detected by the search, i.e. their strain amplitudes are larger than the minimum detectable amplitudes at a given confidence level (the upper limits) of the search. As expected, on average when $50M_{\odot}$ is the maximum BH mass considered, the distance that can be reached is also smaller, since the GW amplitudes scales strongly with the BH mass.
Figure 19: Taken from Abbott et al. (2022d). Using upper limits from the O3 LIGO search for GWs from annihilating boson cloud systems, the maximum distance that the search in Abbott et al. (2022d) could reach at which at least 5% of a simulated BH population would have been detectable, that is, with a strain greater than the value of the upper limit, is shown here. To create this plot, Ref. Abbott et al. (2022d) assumed a uniform spin distribution between [0.2,0.9], different ages, and a Kroupa mass distribution with maximum mass $50M_{\odot}$ (left) and $100M_{\odot}$ (right).
@@ VI.3.2 Cygnus X-1 constraints
Cygnus x-1 is an x-ray binary that is close by, has a large spin, $\sim 0.9$, and has relatively well-measured orbital parameters, which make it a reasonable BH to target. Since the sky position is known, searches based on the Viterbi algorithm discussed in Section VI.2.2 can use much longer coherence times than those used in Abbott et al. (2022d), thus permitting exquisite sensitivity towards almost monochromatic signal, while also being robust to theoretical uncertainties in boson cloud source.
@@
In Fig. 20, we show constraints on the strain amplitude arising from annihilating scalar boson cloud particles as a function of the boson mass. The search results $h_{0}^{95\%}$ are given as the black line, while the two curves represent the numerically obtained strain from this system, accounting for the spin, age, mass and distance of Cyg x-1 ($t_{\rm age}=5\times 10^{6}$ yr and $1\times 10^{5}$ yr). The region in which the search results are below the theoretical curves denote boson masses that are disfavored as having formed around Cyg x-1.
All constraints presented so far do not assume any boson self-interactions. If self interactions do occur, the available parameter space to probe would change. But, if the self-coupling is weak enough, standard searches can probe this scenario without modification. An example for the string axiverse scenario is shown in Fig. 21. The white region indicates the portion of the parameter space in which bosenovas do not occur, and thus cannot be constrained with the current search, while, again, portions of the colored parameter space above the black line indicate that those boson masses are disfavored. However, these constraints contain a number of caveats, namely: (1) the periodic bosenovas do not alter the time-frequency evolution of the GW signal, though they actually might induce oscillations every few minutes Yoshino and Kodama (2015a), and (2) the signal must last for longer than the age of the BH. Future studies are needed to determine to what extent CW searches can actually probe this particular model for string axion boson clouds around BHs, as well as to incorporate self-interactions more systematically into how CW searches are performed.
@@
Figure 20: Taken from Sun et al. (2020), a search using O2 LIGO data for annihilating boson cloud systems around Cyg X-1. (a) Upper limits at 95% confidence on the strain amplitude of annihilating scalar boson clouds coming from the BH in Cyg X-1. The black curve indicates the upper limits from the search on O2 data, with the red stars denoting upper limits derived in frequency bands with injections. The two different curves correspond to different choices for the (unknown) age of the system (lower curve: $10^{6}$ years; upper curve: $5\times 10^{6}$ years. The GW fine structure constant is colored, and the gray region denotes boson masses disfavored under the two different age assumptions. $M=14.8M_{\odot}$, $\chi_{i}=0.99$, and $d=1.86$ kpc for the two curves.
Figure 21: Taken from Collaviti et al. (2025). Constraints on the string axiverse scenario (in blue from O2), in particular the axion decay constant. The regions are divided into three regimes: (1) strong-interactions (above the orange dashed-line), (2) intermediate interactions (between the dashed-blue and dashed orange lines) and no self-interactions (below the dashed-blue line). The three plots correspond to annihilations in different energy levels (left and middle) and to transitions between these energy levels (right). Sensitivities of future detectors are also shown on this plot for comparison.
VI.3.3 Constraints on luminous dark photons clouds

### v4

@@ V.5.4 Ionization
When the binary system separation becomes comparable to the size of the cloud, an “ionization” of the boson cloud occurs as the cloud becomes unbounded from its host BH Takahashi et al. (2022); Baumann et al. (2022a, b). The energy from the binary drives this effect, and is significantly larger than that emitted via GWs; thus, the inspiral of the binary becomes dominated by the boson cloud, instead of simply perturbed by it. Ionization will also tend to circularize the binary if it formed through dynamical capture, while leaving any orbital inclinations unaffected Tomaselli et al. (2023). It has been recently found, however, that as the resonances in the boson cloud evolve during the orbit, the boson cloud is unperturbed and ordinarily visible in LISA if the cloud and binary are counter rotating; otherwise, the cloud is destroyed because of the resonances, but a distinct mark is left on the binary. That is, the binary is forced to co-rotate with the boson cloud, and its eccentricity is driven to a particular value, which allows the possibility of doing statistics with detected binary black hole systems in LISA to probe the boson cloud hypothesis Tomaselli et al. (2024).
Taken together, the various phenomena discussed above—from level transitions and resonances to the dramatic ionization of the cloud—highlight the rich interplay between boson clouds and binary black holes. Each mechanism offers a distinct pathway by which the presence of a boson cloud can alter binary evolution and leave characteristic imprints on the emitted GWs. These theoretical insights provide the foundation for future searches with space-based GW data Bertone et al. (2020); Bertone (2024). In the following section, however, we turn to current ground-based GW interferometer data, which so far constrain only annihilation signals from boson clouds, as discussed in Section V.1.
@@ VI GW constraints on boson clouds
The extensive theoretical background on boson clouds around rotating BHs, coupled with the potential for them to be DM, have motivated GW probes of boson cloud/BH systems. Different GW interferometers probe vastly different mass regimes of ultralight dark matter boson clouds. While ground-based detectors are sensitive to annihilation signals from bosons of masses between $\sim 10^{-14}-10^{-12}$ eV Arvanitaki et al. (2015a), space-based detectors can probe a few orders of magnitude lower than that, $\sim 10^{-20}-10^{-16}$ eV Baumann et al. (2019a), and HFGW detectors could see systems with masses around $\sim 10^{-9}$ eV Aggarwal et al. (2021). At the moment, only ground-based GW interferometers exist, and so we focus on methods and search results from the most recent LIGO, Virgo and KAGRA observing runs. In particular, only GWs from annihilating boson clouds around isolated BHs, discussed in Section V.1, can be searched for and constrained with current GW data.
This section is broken into four parts: in Section VI.1, we will describe potential sources of GWs from annihilating boson clouds that would be visible in LIGO, Virgo and KAGRA. Next, in Section VI.2, we will discuss the CW methods employed to do searches for each of the targets. We will then show recent constraints from different searches for GWs from annihilating boson clouds in Section VI.3, and conclude in Section VI.4 with prospects for future searches for boson clouds with current and future data.
VI.1 Search targets
In Section V, we have explained the different mechanisms that could generate boson clouds around rotating BHs, and thus source GWs through different physical processes – annihilation, transitions, resonances in binary systems, etc. From a practical point of view, we must also discuss where and how we should look for boson clouds. Currently, only ground-based GW interferometers have been built; thus, we are essentially limited to looking for only one of the aforementioned GW emission mechanisms: annihilation, as described in Section V.1. Thus, the next logical question would be: what kind of systems would exhibit annihilation of boson clouds, and where in the sky can they be found?
In principle, there are four targets to consider
  1. 1.
    Known galactic BHs (e.g. Cygnus X-1 Fabian et al. (1989))
  2. 2.
    Known remnants of binary black hole mergers (e.g. GW150914 Abbott et al. (2016))
  3. 3.
    Known binary black hole inspirals and mergers Abac et al. (2025b)
  4. 4.
    Unknown galactic BHs
Each of these sources has its own benefits and drawbacks: known BHs have many of their intrinsic parameters measured – mass, spin, inclination angle, etc. –; however, they are old and thus the strength of GW emission from annihilating boson clouds would be significantly diminished with respect to that when they were first born. On the contrary, remnants of BH mergers are extremely young: their clouds build up quickly and they could begin emitting days or weeks after formation, as described in Section V. Their age is thus precisely measured, as well as the final mass and spin (within $10\%$ in the case of GW150914 Abbott et al. (2016)). These estimates can be even more precise than those that arise from some x-ray binaries. However, remnants of binary black hole mergers are far: GW150914 merged at $0.410$ Mpc, and other detected systems have been seen out to Gpcs. Thus, the scaling of GW amplitude with distance implies that we could not detect remnants of mergers with annihilating scalar boson clouds around them. We note, however, that rapidly annihilating vector boson clouds could be detected out to Gpc Jones et al. (2023, 2025) in current-generation GW interferometers, making merger remnants a viable target even now.
On the other hand, with known binary black hole inspirals, boson clouds could alter the GW signal in a variety of ways, as detailed in Section V.5. Thus, using the spins and masses inferred from LIGO, Virgo and KAGRA measurements, the allowed boson mass can be constrained: if certain BH masses are detected with low spins, this could be evidence that boson clouds have spun down BHs. On the contrary, if BHs with high spins are detected, this could rule out the presence of boson clouds existing in a particular mass range Ng et al. (2021a). Such an analysis allows us to obtain some evidence for or rule out the presence of boson clouds around rotating binary black holes, but does not necessarily permit us to claim confidently that we have detected a boson cloud.
Finally, unknown galactic BHs could take on a wide range of masses, ages and spins depending on when and how they formed. The number of expected BHs in the galaxy is $\mathcal{O}(10^{9})$, which comes from considering the age of the galaxy ($\sim 10^{11}$ years) and the supernova rate ($\sim 1$ per century) Madau and Dickinson (2014). Thus, the BHs we observe are such a small portion of the total number that exist in the local universe, and may not be representative of the full population. However, searches for such systems are entirely blind: we do not know any BH mass or spin parameters, nor do we know the location of such systems. While this may seem to be a daunting task, computationally efficient methods from the CW community have been leveraged to do these kinds of searches D’Antonio et al. (2018); Isi et al. (2019); Jones et al. (2023, 2025).
We will now list the methods that have been developed to search for these sources, and then delve into the constraints that have been set using real LIGO, Virgo and KAGRA data.
@@ VI.2 Search methods
Different methods to search for long-lived GWs from isolated BHs (or those present in a known x-ray binary system) have been developed and used in real searches in LIGO, Virgo and KAGRA data. Detecting GW emission from an isolated BH would provide smoking-gun evidence of the existence of an ultralight bosonic field around BHs. Each has been designed to look for a quasi-monochromatic (scalar boson) or rapidly evolving long-transient (vector boson) signal arising from specific targets, such as Cygnus X-1, or originating from anywhere in the sky. While searches for known systems are significantly less expensive computationally than all-sky ones, they are limited to the few BHs about which we have measurements of distance, spin, mass and age. Thus, a multi-pronged approach has been taken within the community to probe both known and unknown sources of GWs.
VI.2.1 Excess power method
One approach to searching for boson clouds around rotating BHs is the excess-power method, which models the expected signal as approximately monochromatic over the entire observation time. Because the source could lie anywhere on the sky, the signal is modulated by the Earth’s motion, leading to a Doppler-shifted frequency. To recover such a signal, the data are transformed into time–frequency maps and corrected for different sky positions, so that a truly monochromatic signal would appear as a narrow line D’Antonio et al. (2018). This strategy is similar to that described in Section III.1.2, but here the modulation arises from the Doppler effect rather than from the random velocity distribution of individual DM particles. In practice, the analysis is repeated using different coherence times $T_{\text{FFT}}$, which control the width of the frequency bins. Shorter $T_{\text{FFT}}$ values are less sensitive overall but more robust to small frequency drifts, making the search less dependent on specific signal-model assumptions.
In Fig. 15, we show the core stage of this method, which involves creating the time-frequency peakmap from raw time-domain strain data (discussed previously in Section III.1.2), correcting for the Doppler modulation from a particular sky location, and then integrating over time to collect all peaks at a given frequency to highlight the presence of a signal. Note that while we show the color to indicate the ratio of signal to noise power, the method does not actually use it: it only requires that a given point in the time-frequency plane is above a given threshold. When moving from Fig. 15 to Fig. 15, all time-frequency points above this threshold are given a weight of one, and the ones are summed over time in order to create this histogram.
Figure 15: This plot shows how the method described in Section VI.2.1 works for a simulated GW signal from an annihilating scalar boson cloud system. The physics behind this form of DM has been discussed in Section V.1. (a) The Doppler-modulated time-frequency peakmap of the GW signal coming from a particular position in the sky is created by Fourier transforming the GW strain data in chunks of length $T_{\text{FFT}}$, dividing the square of the FFT by the estimation of the power spectral density, and selecting local maxima above a threshold. (b) For a fixed sky position, the time-frequency peakmap is the Doppler corrected, which makes the signal monochromatic. (c) The projected time-frequency map is shown here, which corresponds to an integral over time in each frequency bin. Note that each time-frequency point is labeled simply as “one” if it is above a threshold, and is zero otherwise. This histogram thus is a count of the number of times at which the equalized power exceeded a given threshold.
@@ VI.2.2 Viterbi
Another method that is more robust against stochastic frequency variations relies on Hidden Markov Models (HMMs) to find tracks in the time-frequency plane. The main idea is to model a time-frequency GW signal probabilistically as a Markov chain of transitions between “hidden” (unobservable) frequency states and use a detection statistic to relate “observed” frequency states with the hidden ones Quinn and Hannan (2001). In the framework for this method, the probability to jump to a new state at a given time only depends upon the previous hidden state. In total, there would be $N_{Q}^{N_{T}+1}$ possible paths through the hidden states, where $N_{Q}$ is the total number of hidden states and $N_{T}$ is the total number of times at which we have observed frequency states. The Viterbi algorithm provides a recursive, computationally efficient way of maximizing the probability that the hidden set of states is responsible for the observed sequence of states.
In Fig. 16, taken from Isi et al. (2019), we show the optimal path found by the Viterbi method for injected boson cloud signals in Gaussian noise that has an associated random walk of the frequency, which is meant to simulate unknown theoretical features of the signal. We can see that in both cases, with less and more variability of the signal, the Viterbi algorithm can find the appropriate track.
@@
(a) Less variability
(b) More variability
Figure 16: Taken from Isi et al. (2019). This figure illustrates how the Viterbi HMM algorithm described in Section VI.2.2 works to identify tracks in the time-frequency plane that would arise from GW emission from annihilating boson clouds around rotating BHs. The physics behind this form of DM has been discussed in Section V.1. In particular, these plots show two examples of injected and recovered boson cloud signals with random walk frequency variations $|\delta f|\leq 0.1\Delta f$ (weak variation) and (b) $|\delta f|\leq 0.5\Delta f$ (strong variation), where $\Delta f=1/T_{\rm drift}$ is the frequency bin size for the coherence time of the analysis $T_{\rm drift}=8$ days. The top panels show the injected random walk frequency evolutions, which cannot be seen by eye in the time-frequency tracks. The Viterbi optimal path results in a good match to the injected signal path. Each step in time on the $x$-axis is 8 days. The signal has a frequency of $201.2$ Hz, a spin-up of $10^{-12}$ Hz/s, a randomly chosen sky position, and a strain amplitude of $5\times 10^{-26}$.
@@ VI.2.3 Cross-correlation and other methods
We have already discussed in Section III.1.1 cross correlating detector data to search for ultralight dark matter that could interact directly with GW interferometers; however, it is important to note that such a method could also be used to search for boson clouds around rotating BHs, both deterministically and for a stochastic GW background composed of the superposition of cosmological or astrophysical GW signals. Furthermore, virtually any method in the CW community Riles (2023) could be tuned to look for such boson cloud systems because such methods already analyze almost monochromatic signals focusing on spin-downs but could also extend their parameter spaces to include spin-ups.
VI.3 Search results for scalar boson clouds
Data from the first, second and third observing runs of LIGO, Virgo and KAGRA have been utilized to place constraints on the presence of boson clouds around rotating BHs within our galaxy. One search adopted a general approach, targeting systems across the entire sky. Another focused specifically on GW emission from a boson cloud around Cygnus X-1, and disfavored a range of boson masses. A third search investigated whether luminous dark photon clouds around stellar-mass BHs could mimic known pulsars, setting constraints on the dark-photon mass and its kinetic mixing with the photon. A fourth analysis estimated the stochastic GW backgrounds resulting from the combined effects of scalar, vector, and tensor boson clouds, examining the data for each component individually. Lastly, a fifth study used the masses and spins of observed binary black hole mergers to constrain the BH spin distribution and the mass of the boson. Each of these searches is detailed in the subsections that follow.
@@ VI.3.1 All-sky search constraints
The excess power method described in Section VI.2.1 has been used to analyze LIGO O3 data to search for the presence of annihilating scalar boson clouds around isolated BHs within our galaxy Abbott et al. (2022d, e). This search was very computationally expensive, since each sky position had to be analyzed individually. No evidence for scalar boson cloud systems was found in either case, and thus upper limits were placed on the presence of boson clouds with different ages, spins, and distances away from earth in two different ways Palomba et al. (2019); Dergachev and Papa (2019); Abbott et al. (2022d):
@@
The first way is agnostic towards whether such BHs exist in the galaxy, while the second one contains some (uncertain) astrophysics regarding the mass and spin distributions of astrophysical BHs, but allows for some physical intuition into the reach of such searches.
@@
Figure 17: Taken from Abbott et al. (2022d). Using upper limits from the O3 LIGO search for GWs from annihilating boson cloud systems, exclusion region of particular boson mass /BH mass pairs are derived assuming different ages and distances (left: 1kpc; right: 15 kpc) for BHs with initial spins of $\chi_{i}=0.9$. The physics behind this form of DM has been discussed in Section V.1. This is the most model-agnostic way to probe boson clouds around rotating BHs: the limits do not assume distributions over the mass, spin, distance of BHs in the Galaxy.
Constraints from the second way are shown in Fig. 18, in which a Kroupa BH mass distribution $f(m)\propto M_{\rm BH}^{-2.3}$ Elbert et al. (2018) between $5,~50]M_{\odot}$ and $5,~100]M_{\odot}$ is assumed, as well as a uniform initial BH spin distribution between [0.2,0.9]. For a fixed BH age, the maximum distance reach as a function of boson mass indicates that at least 5% of GWs from BH/boson cloud systems would have been detected by the search, i.e. their strain amplitudes are larger than the minimum detectable amplitudes at a given confidence level (the upper limits) of the search (an example of this quantity is in Section III.1.2, Eq. 71. ). As expected, on average when $50M_{\odot}$ is the maximum BH mass considered, the distance that can be reached is also smaller, since the GW amplitudes scales strongly with the BH mass.
Figure 18: Taken from Abbott et al. (2022d). Using upper limits from the O3 LIGO search for GWs from annihilating boson cloud systems, the maximum distance that the search in Abbott et al. (2022d) could reach at which at least 5% of a simulated BH population would have been detectable, that is, with a strain greater than the value of the upper limit, is shown here. The physics behind this form of DM has been discussed in Section V.1. To create this plot, Ref. Abbott et al. (2022d) assumed a uniform spin distribution between [0.2,0.9], different ages, and a Kroupa mass distribution with maximum mass $50M_{\odot}$ (left) and $100M_{\odot}$ (right). These limits depend on the distributions assumed and are thus less agnostic than those presented in Fig. 17.
@@ VI.3.2 Cygnus X-1 constraints
Cygnus X-1 is an x-ray binary that is close by, has a large spin, $\sim 0.9$, and has relatively well-measured orbital parameters, which make it a reasonable BH to target. Since the sky position is known, searches based on the Viterbi algorithm discussed in Section VI.2.2 can use much longer coherence times than those used in Abbott et al. (2022d), thus permitting exquisite sensitivity towards almost monochromatic signals, while also being robust against theoretical uncertainties in the GW emission from Cygnus X-1, e.g. due to uncertainties on the measured semi-major axis, etc., due to the presence of an accretion disk, etc.
@@
In Fig. 19, we show constraints on the strain amplitude arising from annihilating scalar boson cloud particles as a function of the boson mass using LIGO O2 data. The search results $h_{0}^{95\%}$ are given as the black line, while the two curves represent the numerically obtained strain from this system, accounting for the spin, age, mass and distance of Cygnus X-1 ($t_{\rm age}=5\times 10^{6}$ yr and $1\times 10^{5}$ yr). The region in which the search results are below the theoretical curves denote boson masses that are disfavored as having formed around Cygnus X-1.
All constraints presented so far do not assume any boson self-interactions. If self interactions do occur, the available parameter space to probe would change. But, if the self-coupling is weak enough, standard methods discussed in Section VI.2 can probe this scenario without modification. We show in Fig. 20 constraints on GWs arising from a scalar boson cloud around Cygnus X-1 that have self-interactions. Each panel corresponds to a particular physical mechanism: annihilation in the 211 energy level (left); annihilation in the 322 energy level (middle); and transitions between 322 and 211 energy levels (right). The dashed curves divide the plot into different self-interacting regimes, from top to bottom: strong self-interactions, intermediate self-interactions, and the gravitational regime (negligible self-interactions). The energy level labeling correspond to the quantum numbers $n,l,m$, respectively, where $n$ is the principal quantum number, $l$ is the total angular momentum quantum number, and $m$ is the azimuthal quantum number. The color indicates the expected signal amplitude, which, for different detectors (different colored lines), is mapped to projected constraints on the axion decay constant. The white line indicates the axion decay constant that is fixed by the axion mass.
Constraints on self-interacting DM using the search results on O2 data in Sun et al. (2020) are shown only in the left-hand plot, since that search was designed to analyze 211 annihilation from Cygnus X-1. Furthermore, current and future ground-based interferometers will be able to place constraints on the axion decay constant at frequencies of $\mathcal{O}(100)$ Hz (left plot), while space-based interferometers, such as DECIGO and MAGIS-Space are required to probe transition signals at frequencies below 1 Hz (right panel). The figure also highlights the possibility of multi-band GW astronomy: it would be possible to detect a transition from the 322 to 211 energy levels in space-based observatories while simultaneously observing the CW arising from annihilation of bosons in the 211 energy level in ground-based interferometers Collaviti et al. (2025). Annihilation signals from the 322 level (middle plot) would be too weak to be detected by any of the considered current or future GW interferometers.
@@
Figure 19: Taken from Sun et al. (2020). Upper limits at 95% confidence on the strain amplitude of annihilating scalar boson clouds coming from the BH in Cygnus X-1. The physics behind this form of DM has been discussed in Section V.1. The black curve indicates the upper limits from the search on O2 data, with the red stars denoting upper limits derived in frequency bands with injections. The two different curves correspond to the signal amplitude with different choices for the (unknown) age of the system (lower curve: $10^{6}$ years; upper curve: $5\times 10^{6}$ years. The GW fine structure constant is colored, and the gray regions denote boson masses disfavored under the two different age assumptions. The parameters of Cygnux X-1 used to compute the amplitude curves are: $M=14.8M_{\odot}$, $\chi_{i}=0.99$, and $d=1.86$ kpc.
Figure 20: Taken from Collaviti et al. (2025). Parameter space reachable by current and next-generation GW interferometers for (left) 211 annihilation, (middle) 322 annihilation, and (right) 322 $\rightarrow$ 211 transition signals from a scalar boson cloud around Cygnus X-1, assuming an age of $10^{5}$ yr and the parameters for Cygnus X-1 given in Fig. 19. The numbers 211, 322, and the 322 $\rightarrow$ 211 transition refer to specific energy levels of the cloud. The physics behind this form of DM has been discussed in Sections V.1 and V.2. The dashed curves divide the parameter space into three regimes: (1) top, above orange-dashed line: the strong self-interaction regime, (2) middle, between dashed-orange and dashed-blue lines: the moderate self-interaction regime, and (3) bottom, below the dashed-blue line: the gravitational (negligible self-interactions) regime. The white dotted line shows the expected relation between the axion and its decay constant. Darker colors here correspond to weaker GW signals primarily due to smaller occupation numbers in the energy states of the boson clouds. Constraints on 211 annihilation come from ground-based interferometers, while constraints from transitions will come from space-based interferometers. The 322 annihilation signal is too weak to be detected by any current our future observatory. The O2 constraints come from a real analysis of GW data and are thus representative of both the actual power spectral density of that dataset and an exclusion region on the plot.
VI.3.3 Constraints on luminous dark photon clouds